# MultiBooth: Towards Generating All Your Concepts in an Image from Text

### v2

@@ Abstract
This paper introduces MultiBooth, a method that generates images from texts containing various concepts from users. Despite diffusion models bringing significant advancements for customized text-to-image generation, existing methods often struggle with multi-concept scenarios due to low concept fidelity and high inference cost. MultiBooth addresses these issues by dividing the multi-concept generation process into two phases: a single-concept learning phase and a multi-concept integration phase. During the single-concept learning phase, we employ a multi-modal image encoder and an efficient concept encoding technique to learn a concise and discriminative representation for each concept. In the multi-concept integration phase, we use bounding boxes to define the generation area for each concept within the cross-attention map. This method enables the creation of individual concepts within their specified regions, thereby facilitating the formation of multi-concept images. This strategy not only improves concept fidelity but also reduces additional inference cost. MultiBooth surpasses various baselines in both qualitative and quantitative evaluations, showcasing its superior performance and computational efficiency. Project page is on https://multibooth.github.io/.
@@ 4.1 Comparative Study
Table 3: Quantitative ablation results.
@@
4.2 Ablation Study
Regional Customization Module (RCM). We first verify the effectiveness of RCM by simply removing it. As shown in Fig. 5 and Tab. 3, without RCM, the features of the candle and teapot have fused to some extent. To further validate the effectiveness of RCM, we retain our single concept learning (SCL) and replace our RCM with other layout T2I methods. As shown in the right of Fig. 6, we select two representative methods: LLM-grounded Diffusion (LG) (Lian et al. 2023) and BoxDiff (Xie et al. 2023), with the bounding boxes used displayed on the left. On the one hand, LG (Lian et al. 2023) denoises each concept within the bounding boxes sequentially and then integrates them at the latent level, resulting in concept fusion in the overlapping regions. On the other hand, BoxDiff (Xie et al. 2023) employs the cross-attention map to construct a loss function for updating the latent variables. Although it can generate two concepts simultaneously, it suffers from low image fidelity. Furthermore, neither of these methods can handle complex object interactions according to the given text prompt. In contrast, our method allows different single-concept modules to target specific regions at the cross-attention level, thereby generating multiple concepts simultaneously. By using a base prompt to guide complex object interactions across various regions, we can produce images with both high image fidelity and precise text alignment.
QFormer and Adaptive Concept Normalization (ACN). We also demonstrate the effectiveness of the QFormer and the ACN by removing them either. As shown in Fig. 5 and Tab. 3, without QFormer or ACN, although our method can generate multiple concepts, their fidelity has decreased. In contrast, our full method can faithfully perform multi-concept generation.
@@ 4.3 Discussions
How to handle complex object interaction? Based on Eq. 5, each region is guided by both a regional prompt and a global base prompt. For overlapping regions, the base prompt directly guides their generation along with the regional prompts, ensuring accurate interaction and mitigating concept mixture. As shown in the left of Fig. 6, the absence of the base prompt leads to failures in object interactions or concept feature confusion. However, with the guidance of the base prompt, our method can effectively manage complex object interactions within the overlapping region.
Figure 6: Ablation study on the base prompt and RCM.
@@
Table 4: Inference Time with more concepts.
@@ 2 Concepts 3 Concepts 4 Concepts
4.4 Further Analysis
Customizing more than 2 concepts. Our method theoretically enables the combination of an unlimited number of concepts, facilitating true multi-concept customization. In Fig. 7, we present the results of customizing three and four concepts. The visualization demonstrates that our method consistently generates high-quality customization images when facing a larger number of concepts, further validating the superiority of our approach.
Figure 7: Challenge cases.
Extensions. Our approach is compatible with ControlNet (Zhang, Rao, and Agrawala 2023) for structure-controlled multi-concept generation. Besides, the RCM can guide multiple LoRAs in conducting multi-concept generation and is adaptable to any LoRA-related method, including the LoRA-based DreamBooth (Ruiz et al. 2023). Visualization results are presented in Fig. 8.
Figure 8: Visualization results of extensions.

### v3

@@ Abstract
This paper introduces MultiBooth, a method that generates images from texts containing various concepts from users. Despite diffusion models bringing significant advancements for customized text-to-image generation, existing methods often struggle with multi-concept scenarios due to low concept fidelity and high inference cost. MultiBooth addresses these issues by dividing the multi-concept generation process into two phases: a single-concept learning phase and a multi-concept integration phase. During the single-concept learning phase, we employ a multi-modal image encoder and an efficient concept encoding technique to learn a concise and discriminative representation for each concept. In the multi-concept integration phase, we use bounding boxes to define the generation area for each concept within the cross-attention map. This method enables the creation of individual concepts within their specified regions, thereby facilitating the formation of multi-concept images. This strategy not only improves concept fidelity but also reduces additional inference cost. MultiBooth surpasses various baselines in both qualitative and quantitative evaluations, showcasing its superior performance and computational efficiency.
@@ 4.1 Comparative Study
Figure 5: Qualitative ablation results.
4.2 Ablation Study
Regional Customization Module (RCM). We first verify the effectiveness of RCM by simply removing it. As shown in Fig. 5 and Tab. 3, without RCM, the features of the candle and teapot have fused to some extent. To further validate the effectiveness of RCM, we retain our single concept learning (SCL) and replace our RCM with other layout T2I methods. We select two representative methods: LLM-grounded Diffusion (LG) (Lian et al. 2023) and BoxDiff (Xie et al. 2023), with the bounding boxes used displayed on the left. On the one hand, LG (Lian et al. 2023) denoises each concept within the bounding boxes sequentially and then integrates them at the latent level, resulting in concept fusion in the overlapping regions. On the other hand, BoxDiff (Xie et al. 2023) employs the cross-attention map to construct a loss function for updating the latent variables. Although it can generate two concepts simultaneously, it suffers from low image fidelity. Furthermore, neither of these methods can handle complex object interactions according to the given text prompt. In contrast, our method allows different single-concept modules to target specific regions at the cross-attention level, thereby generating multiple concepts simultaneously. By using a base prompt to guide complex object interactions across various regions, we can produce images with both high image fidelity and precise text alignment.
@@ 4.2 Ablation Study
Table 3: Quantitative ablation results.
QFormer and Adaptive Concept Normalization (ACN). We also demonstrate the effectiveness of the QFormer and the ACN by removing them either. As shown in Fig. 5 and Tab. 3, without QFormer or ACN, the fidelity of our method has decreased. In contrast, our full method can faithfully perform multi-concept generation.
@@ 2 Concepts 3 Concepts 4 Concepts
Table 4: Inference Time with more concepts.